OpenStudy (anonymous):

Please help approach this double integral. I'll type it below... Please help approach this double integral. I'll type it below... @Mathematics

OpenStudy (anonymous):

\[\int\limits_{0}^{2\pi}\int\limits_{0}^{\sqrt{5}}re ^{1+r ^{2}}dr d \theta\]

OpenStudy (anonymous):

haha, after typing the problem I know how to solve it too!

OpenStudy (amistre64):

missing that I see

OpenStudy (amistre64):

multiply by 2/2 and use the top

OpenStudy (amistre64):

missing a 2 ... is what that was spose to say lol \[\int\limits_{0}^{2\pi}\left(\frac{1}{2}\int\limits_{0}^{\sqrt{5}}2r\ e ^{1+r ^{2}}dr\right) d \theta\]